OpenStudy (anonymous):

Alice is selling bracelets to raise money for a local charity. She raise $25 for each box of bracelets that she sells. Her goal is to raise $500 Alice has already sold 3 boxes of bracelets. Write an equation that can be used to determine how many more boxes (x) Alice must sell to meet her goal.

jigglypuff314 (jigglypuff314):

boxes = x 25 each = 25x already sold 3 boxes = (3*25) = 75 goal = 500 (25 each) minus (already sold) equals (goal)